### v1

# Embedded Graph Reconstruction under Hausdorff Noise

    Abstract: Gromov–Hausdorff distance (Def 2) between the abstract graph and a neighborhood graph $\mathcal{N}$ built on the raw data at an appropriate scale, and
  2. 2.
[…]
The reconstruction of metric graphs—both abstract and embedded—has been well-studied in the literature; see for example [2, 6, 7, 8, 10, 11, 12]. Despite the existing reconstruction efforts for abstract graphs under the Gromov–Hausdorff setting, not many successful developments have surfaced to topologically and geometrically reconstruct embedded graphs from Euclidean samples that are only Hausdorff close to the ground truth. For embedded graphs, we particularly mention Majhi [11] for their technique using the Vietoris–Rips complexes. The resulting abstract simplicial complexes remain only topologically faithful, failing to produce a geometric graph as the output in the same ambient space. This paper primarily aims to mitigate such caveats, developing provable reconstruction schemes for embedded graphs from noisy samples under the Hausdorff distance.
Motivation.
We are inspired by the recent pivotal success of Reeb graphs [5, 6, 8] in the reconstruction of abstract metric graphs geometrically. The motivation for this work is two-fold.
First, in the abstract setting, relax the Gromov–Hausdorff assumption to provide reconstruction guarantees from an $(\varepsilon,R)$-approximate (Def 3) sample or neighborhood graph thereof. The Gromov-Hausdorff closeness dictates that the input neighborhood graph is metrically close to the ground truth $G$. In practical applications, this is, indeed, a very strong assumption. A more tenable assumption is that the built neighborhood graph $\mathcal{N}$ and $G$ are metrically close only locally. The notion of $({\varepsilon},R)$-approximation captures the idea.
Secondly, we consider the reconstruction of an embedded metric graph $G\subset{\mathbb{R}}^{d}$ such that the Euclidean distance induces the intrinsic distance $d_{G}$ on $G$ and the raw data $S\subset{\mathbb{R}}^{d}$ is only Hausdorff close to $G$ in ${\mathbb{R}}^{d}$. The latter is typically considered to be a more challenging problem since the Hausdorff closeness does not automatically guarantee a small Gromov–Hausdorff distance. To this end, we put forward the notion of restricted distortion (Def 6) of a graph as a novel sampling parameter for embedded graphs.
Our Contribution.
Under the relaxed $({\varepsilon},R)$-approximated sampling assumptions, we extend the geometric Reeb reconstruction result of [6, 5] for abstract graphs in Theorem 9. Next in Theorem 10, we present our main reconstruction result for embedded graphs. In addition to our theoretical results, we demonstrate our techniques in reconstructing earthquake plate tectonic boundaries from sensor data.
2 Preliminaries
Embedded Metric Graphs
[…]
$d_{G}(x_{1},x_{2})\coloneqq\inf\left\{L(\gamma)\mid\gamma:[a,b]\to G,\text{ joining }x_{1},x_{2}\right\}.$
If $G$ is path-connected, $d_{G}$ defines a metric on $G$.
Metric Spaces
Let $(X,d_{X})$ be a metric space. Let $A$ and $B$ be compact, non-empty subsets. We have the following definition.
Definition 1 (Hausdorff Distance).
The Hausdorff distance between the $A$ and $B$, denoted $d^{X}_{H}(A,B)$, is defined as
$d^{X}_{H}(A,B)\coloneqq\max\left\{\sup_{a\in A}\inf_{b\in B}d_{X}(a,b),\sup_{b\in A}\inf_{a\in B}d_{X}(a,b)\right\}.$
In case $X\subset\mathbb{R}^{N}$ and $A,B,X$ are all equipped with the Euclidean metric, we simply write $d_{H}(A,B)$.
The Gromov–Hausdorff Distance
A correspondence $\mathcal{C}$ between any two metric spaces $(X,d_{X})$ and $(Y,d_{Y})$ is defined as a subset of $X\times Y$ such that (a) for any $x\in X$, there exists $y\in Y$ such that $(x,y)\in\mathcal{C}$, and (b) for any $y\in Y$, there exists $x\in X$ such that $(x,y)\in\mathcal{C}$. We denote the set of all correspondences between $X,Y$ by $\mathcal{C}(X,Y)$. The distortion of a correspondence $\mathcal{C}\in\mathcal{C}(X,Y)$ is defined as:
$\mathrm{dist}(\mathcal{C})\coloneqq\sup_{(x_{1},y_{1}),(x_{2},y_{2})\in\mathcal{C}}|d_{X}(x_{1},x_{2})-d_{Y}(y_{1},y_{2})|.$
Definition 2 (Gromov–Hausdorff Distance).
Let $(X,d_{X})$ and $(Y,d_{Y})$ be two compact metric spaces. The Gromov–Hausdorff distance between $X$ and $Y$, denoted by $d_{GH}(X,Y)$, is defined as:
$d_{GH}(X,Y)\coloneqq\frac{1}{2}\left[\inf_{\mathcal{C}\in\mathcal{C}(X,Y)}\mathrm{dist}(\mathcal{C})\right].$
A slightly weaker and more local notion is $({\varepsilon},R)$- approximation for constants ${\varepsilon},R>0$. The notion has already been used in [2, 9] in the context of reconstruction. However, we use a slightly different version to match the traditional definition of the Gromov–Hausdorff distance.
Definition 3 ($({\varepsilon},R)$-Approximation).
For ${\varepsilon},R$ positive, two metric spaces $(X,d_{X})$ and $(Y,d_{Y})$ are said to be an $({\varepsilon},R)$-approximation of each other if there exists a correspondence $\mathcal{C}\subset X\times Y$ such that for any $(x_{1},y_{1}),(x_{2},y_{2})\in\mathcal{C}$ with
$\min\{d_{X}(x_{1},x_{2}),d_{Y}(y_{1},y_{2})\}\leq R,$
we have $|d_{X}(x_{1},x_{2})-d_{Y}(y_{1},y_{2})|\leq 2{\varepsilon}$.
To see that this is a weaker notion, we note that $d_{GH}(X,Y)<{\varepsilon}$ implies that the two metric spaces are $({\varepsilon},R)$ for any $R>0$. However, an $({\varepsilon},R)$-approximation may not always imply ${\varepsilon}$-closeness in the Gromov–Hausdorff distance; see [9, Example 2.7].
Definition 4 ($\Delta$-Dense).
A path metric space $(X,d_{X})$ is said to be $\Delta$-dense if for $x,x^{\prime}\in X$ there exists a sequence $\{x_{0}=x,x_{1},\cdots,x_{n}=x^{\prime}\}$ such that for all $i=0,\cdots,n-1$, we have $d_{X}(x_{i},x_{i+1})\leq\Delta$ and $d_{X}(x,x^{\prime})=\sum_{i=0}^{n-1}d_{X}(x_{i},x_{i+1})$.
The Reeb Graph
In a metric space $X$, let $d(x)$ be the distance from a fixed base point, and consider the equivalence relation $x\sim y$ if and only if $d(x)=d(y)$ and $x,y$ are in the same path connected component of $d^{-1}(d(x))$. The quotient space $G=X/\sim$ is called the Reeb graph of $d$ and we denote by $\pi:X\rightarrow G$ the quotient map. $\pi$ is continuous and the function $d$ induces a function $d_{*}:G\rightarrow\mathbb{R}_{+}$that satisfies $d=d_{*}\circ\pi$.
The $\alpha$-Reeb Graph
For practical applications, we consider a variant of the Reeb graph that shares properties very similar to the Reeb graph but is more flexible in implementation. Let $\alpha>0$ and let =$\mathcal{I}=\left\{I_{i}\right\}$ be a covering of the range of $d$ by open intervals of length at most $\alpha$. The transitive closure of the relation $x\sim_{\alpha}y$ if and only if $d(x)=d(y)$ and $x,y$ are in the same path connected component of $d^{-1}\left(I_{i}\right)$ for some interval $I_{i}\in\mathcal{I}$ is an equivalence relation that is also denoted by $\sim_{\alpha}$. The quotient space $G_{\alpha}=X/\sim_{\alpha}$ is called the $\alpha$-Reeb graph (associated with the covering $\mathcal{I}$) of $d$ and we denote by $\pi:X\rightarrow G_{\alpha}$ the quotient map. $\pi$ is continuous the function $d$ induces a function $d_{*}:G_{\alpha}\rightarrow\mathbb{R}_{+}$that satisfies $d=d_{*}\circ\pi$.
3 Hausdorff Noise Model
Let $G\subset{\mathbb{R}}^{d}$ be an embedded, compact graph and $S\subset{\mathbb{R}}^{d}$ a finite sample around $G$ such that their Hausdorff distance $d_{H}(S,G)$ is sufficiently small. The sample and the graph come equipped with the standard Euclidean metric.
The Euclidean distance on $G$ induces an intrinsic metric, denoted $d_{G}$, coming from the shortest paths on $G$. In other words, $d_{G}(a,b)$ is the length of the shortest path joining $a,b\in G$.
The Vietoris–Rips Graph
For any compact $Y\subset\mathbb{R}^{d}$, the Vietoris–Rips graph at scale $\eta>0$, denoted $\mathcal{R}_{\eta}(Y)$, is defined as a metric graph in the following manner. Combinatorially, $\mathcal{R}_{\eta}(Y)$ is a graph whose vertex set is $Y$, and a pair $(y_{1},y_{2})\in Y$ forms an edge whenever $\|y_{1}-y_{2}\|\leq\eta$. The combinatorial graph is then metrized in the following way: for each edge $e=(y_{1},y_{2})\in\mathcal{R}_{\eta}(Y)$, take an edge of length $\|y_{1}-y_{2}\|$ that is isometric to the real segment $[0,\|y_{1}-y_{2}\|]$. Then, define a metric on $\mathcal{R}_{\eta}(Y)$ by taking the shortest path metric, which we denote by $d^{\eta}_{Y}$. We note that $\mathcal{R}_{\eta}(Y)$ is $\Delta$-dense for any $\Delta\geq\eta$.
We remark that $d^{\eta}_{Y}(a,b)$ is not finite, in general, for all $\eta>0$. For it to be finite, the Euclidean triangle inequality suggests that the Euclidean thickening $Y^{\frac{\eta}{2}}$ of $Y$ must be path-connected. However, when $d^{\eta}_{Y}$ is finite, $(\mathcal{R}_{\eta}(Y),d^{\eta}_{Y})$ is a metric space as noted in the following proposition. See [9] for a proof.
Proposition 5 (Path-metric Space [9]).
Let $Y\subset{\mathbb{R}}^{d}$ and $\eta>0$ be a number such that the Euclidean thickening $Y^{\frac{\eta}{2}}$ is path-connected. Then, $(\mathcal{R}_{\eta}(Y),d^{\eta}_{Y})$ is a metric space.
Restricted Distortion
Let $G\subset{\mathbb{R}}^{d}$ be an embedded metric graph. The restricted distortion of $G$, denoted $\delta^{\varepsilon}_{R}(G)$, is parametrized by ${\varepsilon}>0$ and $R>0$. Roughly speaking, it is supremum of the ratio of the induced length of metrics on $G$ and (Euclidean) ${\varepsilon}$-thickening of $G$, between points of $G$ that are at least $R$ distance away. More formally, we present the following definition.
Definition 6 (Restricted Distortion).
For $\eta>0$ and $R>0$, the restricted distortion or $(\eta,R)$-distortion of an embedded metric graph $G\subset{\mathbb{R}}^{d}$ is defined as
$\delta^{\eta}_{R}(G)\coloneqq\sup_{d_{G}(x_{1},x_{2})\geq R}\frac{d_{G}(x_{1},x_{2})}{d^{\eta}_{G}(x_{1},x_{2})}.$
It immediately follows from the definition that $\delta^{\eta}_{R}(G)$ is a non-decreasing function of $\eta$ and a non-increasing function of $R$. Moreover, $\delta^{\eta}_{R}(G)\leq\delta(G)$, the global distortion of embedding of $G$. Indeed, we have $\lim_{\begin{subarray}{c}{\varepsilon}\to\infty\\ R\to 0\end{subarray}}\delta^{\varepsilon}_{R}(G)=\delta(G)$. A similar definition has been introduced in [9]. Here, we replace their definition with a more computationally palatable version. As one can easily see, the restricted distortion can be exactly computed for an embedded graph with straight-line edges.
Example 1 (Restricted Distortion of Straight-line Graphs).
Let $G\subset{\mathbb{R}}^{d}$ be an embedded graph with straight-line edges, having $02\theta>$ 0 as the minimum angle between edges. For positive $\eta\leq\frac{R\sin{\theta}}{1-\sin{\theta}}$ and $R$ less than the shortest edge of $G$, we then have
$\delta^{\eta}_{R}(G)=\frac{R\sin{\theta}}{R\sin{\theta}-\eta(1-\sin{\theta})}.$
In case $G$ is compact, a remarkable property of the restricted distortion is that $\delta^{\eta}_{R}(G)\to 1$ uniformly as $\eta\to 0$. See [9] for proof of an alternative version of the restricted version. The same proof should go through in our case.
The uniform convergence provides the following important lemma. The statement has been already proved in [9] for a slightly different restricted distortion and for the raw data $S$, as compared to the Vietoris–Rips graph on $S$. Since $\mathcal{R}_{\eta}(S)$ linearly interpolates the vertices lying on $S$, the proof of [9] applies here without any major modification.
Lemma 7 ($({\varepsilon},R)$–approximation Theorem).
Let $G\subset{\mathbb{R}}^{d}$ be an embedded graph. Let $R>0$ be a fixed number and $\varepsilon\leq\frac{1}{2}R$. If we choose $\eta\in(0,R]$ small enough such that $\delta^{\eta}_{R}(G)\leq\frac{4}{3}$, then for any finite sample $S\subset{\mathbb{R}}^{d}$ with $d_{H}(G,S)<\frac{1}{4}\eta$, the metric space $(\mathcal{R}_{\eta}(S),d^{\eta}_{S})$ is an $\left(\varepsilon,R\right)$-approximation of $(G,d_{G})$.
4 Reconstruction using Reeb Graphs
In this section, we finally prove our main geometric reconstruction result. We begin by showing the following lemma, which generalizes the reconstruction results of [6] under a weaker sampling assumption. A sketch of proof of Lemma 8 is presented in Appendix.
Lemma 8.
Let $(X,d_{X})$ and $(Y,d_{Y})$ be compact geodesic metric spaces, and suppose there are ${\varepsilon},R>0$ such that $X$ is a $R$-dense $({\varepsilon},R)$-approximation of $Y$. Let base points $x_{0}\in X$ and $y_{0}\in Y$ be fixed and let $d_{x_{0}}(x)=d_{X}(x_{0},x)$ and $d_{y_{0}}(y)=d_{Y}(y_{0},y)$ be the distance functions from the base points, and let $M$ be the diameter of $Y$. For $d\geq\alpha\geq 0$, let $D(X,d,\alpha)$ be the maximum diameter of any connected component of $d_{x_{0}}^{-1}([d-\alpha,d+\alpha])$, and let $D(Y,d,\alpha,M,R,{\varepsilon})$ be the maximum diameter of any connected component of $d_{y_{0}}^{-1}([d-\alpha-(\frac{4M}{R}+4){\varepsilon},d+\alpha+(\frac{4M}{R}+4){\varepsilon}])$. Then $D(X,d,\alpha)\leq D(Y,d,\alpha,M,R,{\varepsilon})+(\frac{4M}{R}+2){\varepsilon}$
Lemma 8, along with Proposition 3 of [5], immediately gives the following reconstruction result for $\alpha$-Reeb graphs. For the theorems below, $\beta_{1}(G_{\alpha})$ is the first Betti number of $G_{\alpha}$.
Theorem 9 ($\alpha$-Reeb Reconstruction).
Let $(G,d_{G})$ be a compact metric graph of diameter $M$ and $(X,d_{X})$ a compact geodesic space such that $X$ is $R$-dense and an $({\varepsilon},R)$- approximation of $G$. Let $r\in X$, $\alpha>0$, and $\mathcal{I}$ be a finite covering of the segment $[0,diam{(X)}]$ by open intervals of length at most $\alpha$ such that the $\alpha$-Reeb graph $G_{\alpha}$ associated to $\mathcal{I}$ and the function $d=d_{X}(r,.):X\to\mathbb{R}$ is a finite graph. Then $d_{GH}(X,G_{\alpha})\leq(\beta_{1}(G_{\alpha})+1)[4(2+N_{E,G}(4(\alpha+\frac{4{\varepsilon}M}{R}+4{\varepsilon}))](\alpha+\frac{4{\varepsilon}M}{R}+4{\varepsilon})+(\frac{4M}{R}+2){\varepsilon}$, where $N_{E,G}(\ell)$ is the number of edges of $G$ of length at most $\ell$. Moreover, as $R\to\infty$, this Gromov–Hausdorff approximation reduces to $d_{GH}(X,G_{\alpha})\leq(\beta_{1}(G_{\alpha})+1)[4(2+N_{E,G}(4(\alpha+4{\varepsilon}))](\alpha+4{\varepsilon})+2{\varepsilon}$.
Main Reconstruction Results
We now present our main result to reconstruct the geometry of an embedded graph under the Hausdorff noise.
Theorem 10 (Reconstruction under Hausdorff noise).
Let $G\subset{\mathbb{R}}^{d}$ be a connected, compact embedded graph of diameter $M$. Let $\varepsilon>0$ and $R\in[2\varepsilon,\infty]$. Choose $\eta\in(0,R)$ such that $\delta^{\eta}_{R}(G)\leq\frac{4}{3}$. Let $S\subset{\mathbb{R}}^{d}$ be a compact subset such that $d_{H}(S,G)\\ <\frac{1}{4}\eta$. Let $r\in\mathcal{R}_{\eta}(S)$, $\alpha>0$, and $\mathcal{I}$ be a finite covering of the segment $[0,Diam(\mathcal{R}_{\eta}(S))]$ by open intervals of length at most $\alpha$ such that the $\alpha$-Reeb graph $G_{\alpha}$ associated to $\mathcal{I}$ and the function $d=d_{S}(r,.):\mathcal{R}_{\varepsilon}(S))\to\mathbb{R}$ is a finite graph. Then, $d_{GH}(X,G_{\alpha})\leq(\beta_{1}(G_{\alpha})+1)[4(2+N_{E,G}(4(\alpha+\frac{4{\varepsilon}M}{R}+4{\varepsilon}))](\alpha+\frac{4{\varepsilon}M}{R}+4{\varepsilon})+(\frac{4M}{R}+2){\varepsilon}$, where $N_{E,G}(\ell)$ is the number of edges of $G$ of length at most $\ell$.
Proof.
From Lemma 7, we have that $X=\mathcal{R}_{\eta}(S)$ is an $(\varepsilon,R)$-approximation of $(G,d_{G})$. As we also noted, $X$ is $R$-dense. Theorem 9 then implies that $d_{GH}(X,G_{\alpha})\leq(\beta_{1}(G_{\alpha})+1)[4(2+N_{E,G}(4(\alpha+\frac{4{\varepsilon}M}{R}+4{\varepsilon}))](\alpha+\frac{4{\varepsilon}M}{R}+4{\varepsilon})+(\frac{4M}{R}+2){\varepsilon}$. If the smallest edge length of $G$ is at least $4(\alpha+\frac{4\varepsilon M}{R}+4\varepsilon)$, then the Gromov–Hausdorff distance can be bounded by $(\beta_{1}(G_{\alpha})+1)(8\alpha+\frac{36{\varepsilon}M}{R}+34{\varepsilon})$. ∎
5 Application: Earthquake Plate Tectonic Boundary Reconstruction
Motivated by [6], we seek to apply our techniques of Reeb graph reconstruction to reconstruct plate tectonic boundaries from earthquake epicenter data (Figures 1-5).
Figure 1: Global earthquake occurrences 1970-2009, restricted to magnitude 5 or higher.
Whereas the $\alpha$-Reeb complex is guaranteed to reconstruct the topological structure of the graph with sufficiently small noise, an earthquake fault network is a geospatial object, and the graph reconstruction needs to not only capture the topological structure of the fault network but also follow the location of the faults. As the locations of different earthquakes are measured with differing accuracy in a way that is dependent on their location relative to sensors, a local noise model is appropriate.
Figure 2: $\alpha$-Reeb graph approximation of earthquake data with $\alpha=4$.
Figure 3: Earthquake occurrences 2020-2024 in North America’s intercontinental seismic zone, restricted to magnitude 2.8 or higher.
Figure 4: $\alpha$-Reeb graph approximation of intercontinental seismic zone earthquake data with $\alpha=4$.
Figures 2 and 4 show an example reconstruction of plate tectonic boundaries and the intercontinental seismic zone, respectively, using the $\alpha$-Reeb graph constriction on epicenters of earthquakes (Figures 1 and 3) available in the global earthquake catalog. Currently, the $\alpha$-Reeb construction only uses epicenter locations, measured by latitude and longitude (as in Figure 1). By including features like depth and location error ellipsoids, we predict that our graph reconstruction will be more accurate and useful.
6 Discussions
We present provable reconstruction schemes for graphs embedded in Euclidean space from Hausdorff-close noisy samples. In our approach, we build a neighborhood graph on the sample that is metrically close to the original graph only locally and demonstrate that its $\alpha$-Reeb graph provides a geometrically faithful reconstruction. The success of our geometric reconstruction immediately evokes the question of topological reconstruction of embedded graphs using the same Reeb-based techniques as a future research direction. The nature of our research is ongoing, and we envision making our experimentation with earthquake data more mature and comprehensive. Our reconstruction scheme can benefit several other application areas where the ground truth is an embedded graph.
Appendix A Additional Proofs and Figures
Figure 5: Vietoris–Rips graph constructed from global earthquake data with parameter $r=5$.
Figure 6: Vietoris-Rips graph constructed from earthquake data located in the intercontinental seismic zone in North America with parameter $r=0.75$.
Below we provide a sketch of the proof of Lemma 8
Proof.
Let $\mathcal{C}\subset X\times Y$ be a $({\varepsilon},R)$ correspondence. As $X$ is $R$-dense, distances in $X$ are distorted from corresponding distances in $Y$ by at most $(\frac{4M}{R}+2){\varepsilon}$ (see [2]). Given points $x,x^{\prime}$ and a continuous path $\gamma_{X}$ joining them in a connected component of $d_{x_{0}}^{-1}([d-\alpha,d+\alpha])$, and letting $(x,y),(x^{\prime},y^{\prime})\in\mathcal{C}$, we can construct (by subdividing the path $\gamma_{X}$ in $R$-short subpaths) a continuous path $\gamma_{Y}$ joining $y,y^{\prime}$ in a connected component of $d_{y_{0}}^{-1}([d-\alpha-(\frac{4M}{R}+4){\varepsilon},d+\alpha+(\frac{4M}{R}+4){\varepsilon}])$. The result then follows.